OpenStudy (anonymous):

Write a formula for f of g of h: f(x) = sq. root x+1 g(x) = 1/(x+4) h(x) = 1/x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

h(X)=1/x g(h(x))=1/(1/x+4)) g(h(x))=1/(1+4x) f(g(h(x)))=sqrt(1/(1+4x)+1) and simplify